OpenStudy (anonymous):

What is the domain of the graph of y = −2x2 − 4x + 7?

OpenStudy (anonymous):

pleasee help me

OpenStudy (anonymous):

-2x^2 - 4x + 7 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x can be all real numbers. If you make x be 2 then the answer would be -2(2)^2 - 4(2) + 7 which equals -9 which is still a real number.

OpenStudy (anonymous):

thank youu

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

can you help me in another one?

OpenStudy (anonymous):

Where does the graph of y = 6x2 − 5x − 6 cross the x-axis? (−, 0) and (6, 0) (, 0) and (−6, 0) (, 0) and (−, 0) (−, 0) and (, 0)

OpenStudy (anonymous):

well first you need to factor them down into the form (2x + a) (3x +b) and then set that equal to zero and solve for x to get your answer.

OpenStudy (anonymous):

can u help me do that?

OpenStudy (anonymous):

I'm working on it.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The factored form of that is (2x - 3) (3x + 2)

OpenStudy (anonymous):

If you set that to zero you get two equations 2x - 3 = 0 3x + 2 = 0 Solving for x we get x = 3/2 and x = -2/3

OpenStudy (anonymous):

yes! okay now what?

OpenStudy (anonymous):

Well that's the wrong answer.... first of all. erm...

OpenStudy (anonymous):

Are you sure you have the right question? because 3/2 and -2/3 should be the answer.

OpenStudy (anonymous):

yes thats what i got!

OpenStudy (anonymous):

I'm sorry I don't know.